OpenStudy (anonymous):

Solve: 8x - 10 + 18 = 24 Answer -2 4 6.5 2 (= Thanks.

OpenStudy (anonymous):

8x + 8 = 24 8x = 16 x = 2

OpenStudy (lilg132):

bring all numbers to one side and have all variables on the other side so you would get 8x = 24 + 10 - 18 8x = ? then just divide both sides by the number next to x which will solve x for you which is 8 dont forget to click good answer :)

OpenStudy (lilg132):

moneybird you give the answer away too easily

OpenStudy (anonymous):

I show steps too

OpenStudy (lilg132):

try and get them to give you the answer, just a suggestion